# Applications of Automatic Differentiation in Image Registration

### v1

@@ 4.3 Implementation and Results
Figure 4: Our test problem for super resolution. The reference image is $20\times 20$ pixels, and the $4$ templates are $10\times 10$ pixels. The problem is well-determined, as we are to infer $400$ parameters from $400$ pixels.
We consider the small super-resolution problem shown in Figure 4 and note a significant benefit in reconstruction quality and convergence rate of the Gauss-Newton method when using AD to differentiate through all iterations of CG (see Figure 5 and Figure 6). There are $18$ registration parameters to estimate in this problem, and we found that computing the Jacobian of the variable-projected objective function (Equation 7) using forward-mode automatic differentiation takes approximately six to ten times longer than evaluating the objective function itself, which is in line with our expectations. The final relative reconstruction error when differentiating all iterations of CG is $10.7\%$, while ignoring the Jacobian of the projection results in a higher relative reconstruction error of $25.61\%$.
At each optimization step, the Jacobian 8 is computed using forward mode AD. Its important to note that in forward mode, memory consumption scales with the size of the input rather than the number of operations in a forward pass. For the variable projected super-resolution problem, the memory consumption due to forward mode should remain constant regardless the number of CG iterations. In our experiments we observed that memory usage increases with the number of optimization steps which can be attributed to an implementation issue in PyTorch.
@@ 4.3.1 Differentiation of a Portion of CG Iterations
Figure 7: Jacobian convergence plot. $\text{Error}=\|f(x+hv)-f(x)-h\nabla f(x)\|$ where $h$ is the perturbation multiplier, $v$ is the direction of the perturbation, and $\nabla f(x)$ varies depending on the percentage of final iterations of CG that are differentiated.
Differentiating through the later iterations of CG proves essential for accurate computation. In these experiments, a total of $200$ iterations of CG are used in each optimization step, but the percentage of the final iterations that are differentiated through vary. Figure 7 presents convergence plots for Jacobians obtained via AD, comparing various percentages of CG iterations through which differentiation is applied. Interestingly, differentiating through the final 70% of iterations leads to poorer convergence than not differentiating through any iterations at all. However, when differentiating through the entire set of iterations (100%), the convergence rate surpasses that of the 90% and 95% cases.
@@
Our results shown in Figure 8 demonstrate that differentiating through more than $70$% is necessary. Differentiating through $100$% of the iterations of CG is the best in terms of relative loss and gradient norm reduction. The final reconstruction qualities additionally indicate that $100$% is the best, but the visual difference in error is small after $90$% as seen in Figure 9. We also tested differentiating through the initial iterations of CG, but the performance was worse than using the later iterations. Therefore, these results are not included in this work.
@@ 4.3.2 Using Inexact Projections
Figure 10: The performance of $10$ iterations of variable-projected Gauss-Newton on the problem from Figure 4 for various numbers of Conjugate Gradient Iterations used in each projection. The final relative error is computed by solving the least squares problem exactly for the final reference image prediction and measuring the difference from the true reference used to create the data.
Our results (Figure 10) show that there are diminishing returns for using more exact projections in each iteration. Using $50$ or $100$ iterations performed the best in terms of relative loss and gradient norm reduction. The final recovered reference images (Figure 11) show that the best reconstruction qualitatively is that using $100$ CG iterations, but $10$, $25$, and $50$ are close to the same reconstruction quality and the visual difference in reconstructions and error images is small.
@@ 5 Conclusions
In the problem of super resolution, we showed that computing the full Jacobian of a variable projected objective function improves both the speed of convergence of the optimization and the final reconstruction quality from the previous approach of not differentiating through the projection, though it introduces computational overhead. Through experiments on a small super resolution problem, we found that tracking the computations of at least $70$% of the final iterations of the CG method was necessary to see the benefits of including the Jacobian of the projection. Moreover, we found that using inexact projections during each iteration of optimization maintains high reconstruction quality with fewer CG iterations, providing a balance between computational efficiency and accuracy. These results suggest that leveraging inexact projections can mitigate computational costs while maintaining the advantage in results that the AD computed Jacobian of the projection provide.

### v2

@@ 4.3 Implementation and Results
Figure 4: Our test problem for super resolution. The reference image is $\times 20$ 20 pixels, and the 444 templates are $\times 10$ 10 pixels. The problem is well-determined, as we are to infer 400400400 parameters from 400400400 pixels.
[…]
At each optimization step, the Jacobian (8) is computed using forward mode AD. Its important to note that in forward mode, memory consumption scales with the size of the input rather than the number of operations in a forward pass. For the variable projected super-resolution problem, the memory consumption due to forward mode should remain constant regardless the number of CG iterations. In our experiments we observed that memory usage increases with the number of optimization steps which can be attributed to an implementation issue in PyTorch.
@@ 4.3.1 Differentiation of a Portion of CG Iterations
Figure 7: Jacobian convergence plot for the residual function ${\bf r}({\bf w})=\mathsf{A}_{\lambda_{{\bf f}}}({\bf w}){\bf f}^{(0)}({\bf w})-{\bf b}$. In this case, $\text{Error}=\|{\bf r}({\bf w}+h{\bf v})-{\bf r}({\bf w})-h\frac{\partial bfr({\bf w})}{\partial{\bf w}}{\bf v}\|,$ where $h$ is the perturbation multiplier, $v$ is the direction of the perturbation, and our Jacobian approximation $\frac{\partial bfr({\bf w})}{\partial{\bf w}}$ varies in accuracy depending on the percentage of final iterations of CG that are differentiated.
Differentiating through the later iterations of CG proves essential for accurate computation. In these experiments, a total of 200200200 iterations of CG are used in each optimization step, but the percentage of the final iterations that are differentiated through vary. Figure 7 presents convergence plots for Jacobians obtained via AD, comparing various percentages of CG iterations through which differentiation is applied. Interestingly, differentiating through the final 70% of iterations leads to poorer convergence than not differentiating through any iterations at all. However, when differentiating through the entire set of iterations (100%), the convergence rate surpasses that of the 90% and 95% cases.
@@
Our results shown in Figure 8 demonstrate that differentiating through more than 707070% is necessary. Differentiating through 100100100% of the iterations of CG is the best in terms of relative loss and gradient norm reduction. The final reconstruction qualities additionally indicate that 100100100% is the best, but the visual difference in error is small after 909090% as seen in Figure 9. We also tested differentiating through the initial iterations of CG, but the performance was worse than using the later iterations. Therefore, these results are not included in this work.
@@ 4.3.2 Using Inexact Projections
Figure 10: The performance of 101010 iterations of variable-projected Gauss-Newton on the problem from Figure 4 for various numbers of Conjugate Gradient Iterations used in each projection. The final relative error is computed by solving the least squares problem exactly for the final reference image prediction and measuring the difference from the true reference used to create the data.
Our results (Figure 10) show that there are diminishing returns for using more exact projections in each iteration. Using 505050 or 100100100 iterations performed the best in terms of relative loss and gradient norm reduction. The final recovered reference images (Figure 11) show that the best reconstruction qualitatively is that using 100100100 CG iterations, but 101010, 252525, and 505050 are close to the same reconstruction quality and the visual difference in reconstructions and error images is small.
@@ 5 Conclusions
In the problem of super resolution, we showed that computing the full Jacobian of a variable projected objective function improves both the speed of convergence of the optimization and the final reconstruction quality from the previous approach of not differentiating through the projection, though it introduces computational overhead. Through experiments on a small super resolution problem, we found that tracking the computations of at least 707070% of the final iterations of the CG method was necessary to see the benefits of including the Jacobian of the projection. Moreover, we found that using inexact projections during each iteration of optimization maintains high reconstruction quality with fewer CG iterations, providing a balance between computational efficiency and accuracy. These results suggest that leveraging inexact projections can mitigate computational costs while maintaining the advantage in results that the AD computed Jacobian of the projection provide.